OpenStudy (chacha1234):

Can someone please help me with this quick question? it's the last one on my 8th grade homework Sylvain spent a total of 18 hours on English and math homework over the last 3 weeks. He spent twice as much time on English than on math. Let x represent the number of hours spent on English and y represent the number of hours spent on math. x+y=18 x=2y What is the solution of the system of equations? (12, 6) (6, 12) (9, 9) (15, 3)

OpenStudy (xguardians):

Substitute: 2y + y = 18 3y = 18 y = 6 Resub: x = 2y x = 2(6) x = 12 (12,6)

OpenStudy (chacha1234):

@xGuardians Aisha bought 15 apples and oranges for $9.00. Each orange costs $0.50, and each apple costs $0.65. Let x represent the number of oranges and y represent the number of apples. Which system can be used to find the number of apples and oranges Aisha bought?

OpenStudy (chacha1234):

Why did you change your profile pic?

OpenStudy (xguardians):

.5x + .65y = $9.00 x+y = 15 Solve for one side x = 15 - y Resub .5(15-y)+.65y = 9.00 7.5 - .5y + .65y = 9.00 -.5y + .65y = 1.5 .15y = 1.5 y = 10 Resub again x+y=15 x + 10 = 15 x = 5 Orange = x Apple = y Orange = 5 Apple = 10